# Spiking NeRF: Representing the Real-World Geometry by a Discontinuous Representation

### v2

@@ Introduction
One of the reasons for the success of NeRF is its ability to implicitly represent geometric information using a neural network based on conventional MLPs, which are continuous functions. However, geometric information in the real world is discontinuous at the interface between the air and the surface, which is inconsistent with the computational representation in conventional MLPs. This inconsistency poses three problems. Post-processing. Existing methods require a post-processing approach (i.e., filtering with a threshold (Peng et al. 2021; Pumarola et al. 2021; Boss et al. 2021)) to extract discontinuous geometric representations from learned continuous fields. However, determining the optimal threshold requires empirical and tedious tuning, and many methods rely on manual selection of the threshold for different scenarios (Mildenhall et al. 2021; Pumarola et al. 2021; Boss et al. 2021) (see the left of Fig. 1). Optimal threshold perturbation. NeRF-based methods optimize the network view by view, resulting in the optimal threshold of the surface to perturbate (see supplementary material for more details). However, existing methods use the same threshold to filter output values (Mildenhall et al. 2021; Peng et al. 2021; Pumarola et al. 2021; Boss et al. 2021), which cannot obtain accurate geometric information because the optimal filtering thresholds under different views vary (see the middle of Fig. 1). Light density scenarios [^1] These scenarios contain several rays, along which the density with non-zero values distribute within a very narrow range (e.g., thin objects) or with small values (e.g., semi-transparent objects). In real life, semi-transparent objects (e.g. windows, glasses) and thin objects (e.g. branches, axles, nets) can be seen everywhere.. The inconsistency can result in even greater errors in light density scenarios (see the right of Fig. 1). Because it is difficult for continuous functions like conventional MLPs to generate quite different densities (i.e., significantly different outputs) for points close to each other (i.e., similar inputs), resulting in the disappearance of the object.
@@
Although there are some attempts to address the aforementioned problems, these methods are still based on continuous representation and do not provide a fundamental solution to those problems. For example, some works replace the density field with other fields to avoid threshold selection (e.g., (Oechsle, Peng, and Geiger 2021; Wang et al. 2021; Yariv et al. 2021; Wang, Skorokhodov, and Wonka 2022)). However, these methods cannot use the filter value to get an accurate geometric representation (see the left of Fig. 1). Moreover, they fail on light density scenarios (see the right of Fig. 1). Some works modify the input to improve its frequency or discretization, thereby mitigating the inconsistency between continuous and discontinuous representations. However, these methods still fail on light density scenarios (Wang, Skorokhodov, and Wonka 2022, 2023) (see the right of Fig. 1). Some works are designed for specific scenarios (e.g., (Guo et al. 2022; Levy et al. 2023)). However, these methods have poorer performance in general scenarios. Moreover, they mainly focus on novel view synthesis, and the quality of the 3D reconstruction is low (see the right of Fig. 1).
@@
Recently, neural implicit functions have emerged as an effective representation of 3D geometry (Atzmon et al. 2019; Peng et al. 2020; Saito et al. 2019; Xu et al. 2019) and appearance (Mildenhall et al. 2021; Liu et al. 2020a, b; Schwarz et al. 2020; Sitzmann, Zollhöfer, and Wetzstein 2019) as they represent 3D content continuously and without discretization while simultaneously having a small memory footprint. Most of these methods require 3D supervision. However, several recent works (Mildenhall et al. 2021; Niemeyer et al. 2020; Yariv et al. 2020) demonstrated differentiable rendering for training directly from images. Some works use estimated depth information for surface rendering without pixel-accurate object masks (Verbin et al. 2022). Some works enhance the accuracy of geometric information by incorporating point cloud information and warp operations (Fu et al. 2022; Darmon et al. 2022). Some works do not model the density field and use other fields to avoid threshold selection (Oechsle, Peng, and Geiger 2021; Wang et al. 2021; Fu et al. 2022; Darmon et al. 2022; Yariv et al. 2021; Wang, Skorokhodov, and Wonka 2022). However, these works are still based on a continuous representation, the same threshold cannot accurately divide the surface. Different from previous methods, we model the 3D geometric information in a discontinuous representation by proposing a hybrid ANN-SNN framework.
@@
Over the past few years, brain-inspired (Li et al. 2021a) deep SNNs using spike trains (Marchisio et al. 2020) have gradually replaced ANNs in various tasks (Zhang, Huang, and He 2023; Zhang et al. 2023; Ororbia 2023), and it is extensively utilized to develop energy-efficient solutions for various tasks. Some works use SNNs to process video or event streams (Zhu et al. 2022; Gehrig et al. 2020). Some works exploit the discontinuity of the network to enhance its robustness (Bagheri, Simeone, and Rajendran 2018; Sharmin et al. 2020). Some works leverage the neural parameter of SNNs to further improve the utilization of available information (Li et al. 2022; Zhang et al. 2022). Although previous works have demonstrated SNN applications on a wide range of tasks, they are still limited in their performance (Neftci, Mostafa, and Zenke 2019; Li et al. 2021b; Fang et al. 2021). Meanwhile, there has been a growing interest in exploring the potential benefits of combining ANNs and SNNs (Kugele et al. 2021; Lee et al. 2020; Liu and Zhao 2022; Zhao et al. 2022). By combining ANNs and SNNs, better performance can be achieved while reducing the time step of SNNs. Different combination strategies have been explored for a variety of tasks. A group of work employs the strategy of processing the accumulated spike train of SNNs with ANNs (Kugele et al. 2021; Lee et al. 2020; Liu and Zhao 2022). In these works, the SNN is used as an efficient encoder of spatio-temporal data. The output of the SNN is accumulated to summarize the temporal dimension before the ANN processes the accumulated features (Kugele et al. 2021; Lee et al. 2020). A second line of work uses a strategy where the output of the independently operating SNN and ANN is fused (Lele et al. 2022; Zhao et al. 2022). In these works, the outputs of both networks are fused based on heuristics (Lele et al. 2022) or accumulation based on the output of the ANN (Zhao et al. 2022). Different from the previous methods, we apply the hybrid ANN-SNN framework to a non-cognitive application.
@@
Here, p refers to each pixel as defined in (Guo et al. 2022). There are currently two main methods for depth estimation: one based on the integration (Zhang et al. 2021; Deng et al. 2022) and the other based on the threshold to get global shape (Wang et al. 2021; Ichnowski et al. 2021; Fu et al. 2022). Since we focus on a global geometric shape extracted from the network by threshold rather than a depth map dependent on perspective, this paper mainly follows the second method to define the depth.
@@
We introduce the well-known Integrate-and-Fire (IF) neuron model (Leroy, Franco, and Boyer 2018) for SNNs. Given a membrane potential $u_{t}$ at time step $t$, the membrane potential $u^{\text{pre}}_{t+1}$ before firing at time step ${t+1}$ is updated as:
@@
Here, the elements, $W$ and $x_{t+1}$, represent respectively the weight and the output from the previous layer at time step ${t+1}$. ${V_{\text{th}}}$ is the firing threshold. The spiking neuron will fire a spike when $u^{\text{pre}}_{t+1}$ exceeds the ${V_{\text{th}}}$ (see Fig. 2), and then the membrane potential $u_{t+1}$ at time step ${t+1}$ resets to 00. The spike output is given by:
@@
However, the spikes convey less information than the floating-point number, and networks using IF neurons do not perform well in regression tasks (Gehrig et al. 2020; Rançon et al. 2022; Lu et al. 2021; Shrestha and Orchard 2018). It’s difficult to ensure accuracy by directly using IF neurons for modeling. To obtain full-precision information, (Li et al. 2022) changes the Eqn. 4 to:
@@
B-FIF: Bounded full-precision integrate and fire spiking neuron.
@@
$u^{\text{pre}}=k\cdot r\cdot\tanh({u+Wx\over r}),$ (9)
@@
${{\partial y}\over{\partial V_{\text{th}}}}=\lambda\max(0,{k-\left|u^{\text{pre}}-V_{\text{th}}\right|\over k^{2}}).$ (14)
@@
We evaluate our methods in 8 scenes from the Blender dataset (Mildenhall et al. 2021) and 6 scenes from Dex-NeRF (Ichnowski et al. 2021). Following most 3D reconstruction methods (Wang et al. 2021; Wang, Skorokhodov, and Wonka 2022, 2023; Oechsle, Peng, and Geiger 2021), we also evaluate our methods in 15 scenes from the DTU dataset (Jensen et al. 2014). These datasets cover different types of scenes or objects and are benchmarks for NeRF-based methods. Furthermore, we additionally use 2 self-created scenes to show that our method is better than previous methods in special scenarios (see supplementary material for more details).
@@
Method wheel tennis mount glass turbine clutter pawn pipe Avg.
@@
As shown in the 5th and 6th columns of Fig. 3, with conventional NeRF-based methods, determining the threshold that can accurately segment the surface requires enumeration attempts. We choose the optimal threshold manually for them. Due to the lack of surface constraints, the reconstruction results have many erroneous high-frequency information, and there is a loss of reconstruction accuracy for thin objects and semi-transparent scenes. As shown in the 2nd, 3rd, and 4th columns of Fig. 3, with SDF-based methods, the reconstruction resulted in significant errors in high-frequency changes in the image, such as failing to reconstruct holes that should have been present and surfaces that should have been separate. Moreover, NeuS demonstrated excessive smoothness in the 2nd row of Fig. 3. NeuS, HF-NeuS, and PET-NeuS fail in the reconstruction of semi-transparent scenes. Even though HF-NeuS uses some methods to reconstruct the high-frequency information, it still fails in some thin high-frequency positions. Because our method uses the hybrid ANN-SNN framework to model the 3D geometric information in a discontinuous representation, the reconstructed geometric information is more accurate.
@@
Method lego chair mic ficus hotdog drums materials ship Avg.
PET 0.58 0.65 0.59 0.71 1.02 2.53 1.05 1.57 1.09
HFS 0.96 0.65 0.72 0.87 1.35 3.82 1.08 2.18 1.45
NeuS 1.52 0.70 0.85 1.67 1.40 4.27 1.08 2.33 1.73
NeRF 2.06 0.75 0.95 0.56 1.83 3.38 1.12 3.12 1.72
Mip 1.92 0.90 1.13 0.55 1.98 3.34 1.30 2.03 1.64
Ours 0.70 0.66 0.72 0.54 0.94 2.43 1.10 1.49 1.07
Table 3: Quantitative Comparison on DTU (Jensen et al. 2014). We show a quantitative comparison for the reconstructed geometry on 15 scans from the DTU dataset.
Method 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Avg.
PET 0.59 0.84 0.72 0.40 0.95 0.84 0.72 1.42 1.14 0.72 0.53 1.05 0.44 0.59 0.68 0.78
HFS 0.83 1.37 0.78 0.47 1.11 0.68 0.68 1.20 1.17 0.85 0.57 1.27 0.38 0.54 0.55 0.83
NeuS 1.00 1.37 0.93 0.43 1.10 0.70 0.72 1.48 1.16 0.83 0.52 1.69 0.35 0.49 0.54 0.89
NeRF 1.90 1.63 1.75 0.60 2.03 1.07 1.47 1.70 1.95 1.13 0.79 2.33 0.87 1.05 0.86 1.41
Mip 2.15 1.83 1.65 1.61 2.90 1.79 1.52 1.93 2.19 1.32 1.79 2.52 1.38 1.49 1.21 1.82
Ours 0.84 1.20 1.02 0.38 1.15 0.72 0.69 1.10 1.19 0.65 0.49 1.60 0.49 0.55 0.51 0.83
@@
To assess the effectiveness of the initial smoothing, we design an experiment without $L_{\text{g}}$ and compare the result with our method. It can be seen that although the Chamfer distances are similar, the reconstructed surface contains wrong high-frequency information (see the 4th column of Fig. 5), indicating that training from the wrong high-frequency surface is difficult and training from a smooth surface is necessary.
@@ Conclusion
We present a novel neural surface reconstruction method for reconstructing objects with high fidelity from 2D image inputs. NeRF and its variants use volume rendering to produce a neural scene representation. However, extracting high-quality surfaces from this learned implicit representation is difficult because there is not a definite interface density. In our work, we propose to use a hybrid ANN-SNN framework to reconstruct density fields. We observe that the conventional volume rendering method causes inherent geometric errors for surface reconstruction, and propose a new spiking neuron to get more accurate surface reconstruction with an adaptive spiking threshold.

### v3

@@ Introduction
One of the reasons for the success of NeRF is its ability to implicitly represent geometric information using a neural network based on conventional MLPs, which are continuous functions. However, geometric information in the real world is discontinuous at the interface between the air and the surface, which is inconsistent with the computational representation in conventional MLPs. This inconsistency poses three problems. Post-processing. Existing methods require a post-processing approach (i.e., filtering with a threshold (Peng et al. 2021; Pumarola et al. 2021; Boss et al. 2021)) to extract discontinuous geometric representations from learned continuous fields. However, determining the optimal threshold requires empirical and tedious tuning, and many methods rely on manual selection of the threshold for different scenarios (Mildenhall et al. 2021; Pumarola et al. 2021; Boss et al. 2021) (see the left of Fig. 1). Optimal threshold perturbation. NeRF-based methods optimize the network view by view, resulting in the optimal threshold of the surface to perturbate (see supplementary material for more details). However, existing methods use the same threshold to filter output values (Mildenhall et al. 2021; Peng et al. 2021; Pumarola et al. 2021; Boss et al. 2021), which cannot obtain accurate geometric information because the optimal filtering thresholds under different views vary (see the middle of Fig. 1). Light density scenarios [^1] These scenarios contain several rays, along which the density with non-zero values distribute within a very narrow range (e.g., thin objects) or with small values (e.g., semi-transparent objects). In real life, semi-transparent objects (e.g., windows, glasses) and thin objects (e.g., branches, axles, nets) can be seen everywhere.. The inconsistency can result in even greater errors in light density scenarios (see the right of Fig. 1). Because it is difficult for continuous functions like conventional MLPs to generate quite different densities (i.e., significantly different outputs) for points close to each other (i.e., similar inputs), resulting in the disappearance of the object.
@@
Although there are some attempts to address the aforementioned problems, these methods are still based on continuous representation and do not provide a fundamental solution to those problems. For example, some works replace the density field with other fields to avoid threshold selection (e.g., (Oechsle, Peng, and Geiger 2021; Wang et al. 2021; Yariv et al. 2021)). However, these methods cannot use the filter value to get an accurate geometric representation (see the left of Fig. 1). Moreover, they fail on light density scenarios (see the right of Fig. 1). Some works modify the input to improve its frequency or discretization, thereby mitigating the inconsistency between continuous and discontinuous representations. However, these methods still fail on light density scenarios (Wang, Skorokhodov, and Wonka 2022, 2023) (see the right of Fig. 1). Some works are designed for specific scenarios (e.g., (Guo et al. 2022; Levy et al. 2023)). However, these methods have poorer performance in general scenarios. Moreover, they mainly focus on novel view synthesis, and the quality of the 3D reconstruction is low (see the right of Fig. 1).
@@
Recently, neural implicit functions have emerged as an effective representation of 3D geometry (Atzmon et al. 2019; Peng et al. 2020; Saito et al. 2019; Xu et al. 2019) and appearance (Liu et al. 2020a, b; Schwarz et al. 2020; Sitzmann, Zollhöfer, and Wetzstein 2019) as they represent 3D content continuously and without discretization while simultaneously having a small memory footprint. Most of these methods require 3D supervision. However, several recent works (Niemeyer et al. 2020; Yariv et al. 2020) demonstrated differentiable rendering for training directly from images. Some works use estimated depth information for surface rendering without pixel-accurate object masks (Verbin et al. 2022). Some works enhance the accuracy of geometric information by incorporating point cloud information and warp operations (Fu et al. 2022; Darmon et al. 2022). Some works do not model the density field and use other fields to avoid threshold selection (Oechsle, Peng, and Geiger 2021; Wang et al. 2021; Fu et al. 2022; Darmon et al. 2022; Yariv et al. 2021). However, these works are still based on a continuous representation, the same threshold cannot accurately divide the surface. Different from previous methods, we model the 3D geometric information in a discontinuous representation by proposing a hybrid ANN-SNN framework.
@@
Over the past few years, brain-inspired (Li et al. 2021a) deep SNNs using spike trains (Marchisio et al. 2020) have gradually replaced ANNs in various tasks (Zhang, Huang, and He 2023; Zhang et al. 2023; Ororbia 2023), and it is extensively utilized to develop energy-efficient solutions for various tasks. Some works use SNNs to process video or event streams (Zhu et al. 2022; Gehrig et al. 2020). Some works exploit the discontinuity of the network to enhance its robustness (Bagheri, Simeone, and Rajendran 2018; Sharmin et al. 2020). Some works leverage the neural parameter of SNNs to further improve the utilization of available information (Li et al. 2022; Zhang et al. 2022). Although previous works have demonstrated SNN applications on a wide range of tasks, they are still limited in their performance (Li et al. 2021b; Fang et al. 2021). Meanwhile, there has been a growing interest in exploring the potential benefits of combining ANNs and SNNs (Kugele et al. 2021; Liu and Zhao 2022). By combining ANNs and SNNs, better performance can be achieved while reducing the time step of SNNs. Different combination strategies have been explored for a variety of tasks. A group of work employs the strategy of processing the accumulated spike train of SNNs with ANNs (Kugele et al. 2021; Lee et al. 2020). In these works, the SNN is used as an efficient encoder of spatio-temporal data. The output of the SNN is accumulated to summarize the temporal dimension before the ANN processes the accumulated features (Kugele et al. 2021; Lee et al. 2020). A second line of work uses a strategy where the output of the independently operating SNN and ANN is fused (Lele et al. 2022; Zhao et al. 2022). In these works, the outputs of both networks are fused based on heuristics (Lele et al. 2022) or accumulation based on the output of the ANN (Zhao et al. 2022). Different from the previous methods, we apply the hybrid ANN-SNN framework to a non-cognitive application.
@@
Here, p refers to each pixel. There are currently two main methods for depth estimation: one based on the integration (Zhang et al. 2021; Deng et al. 2022) and the other based on the threshold to get global shape (Wang et al. 2021; Ichnowski et al. 2021). Since we focus on a global geometric shape extracted from the network by threshold rather than a depth map dependent on perspective, this paper mainly follows the second method to define the depth.
@@
We introduce the well-known Integrate-and-Fire (IF) model (Dayan and Abbott 2005). Given a membrane potential $u_{t}$ at time step $t$, the membrane potential $u^{\text{pre}}_{t+1}$ before firing at time step ${t+1}$ is updated as:
@@
Here, $W$ and $x_{t+1}$ represent respectively the weight and the output from the previous layer at time step ${t+1}$. ${V_{\text{th}}}$ is the firing threshold. The spiking neuron will fire a spike when $u^{\text{pre}}_{t+1}$ exceeds ${V_{\text{th}}}$ (see Fig. 2), and then $u_{t+1}$ resets to 00. The spike output is given by:
@@
However, the spikes convey less information than the floating-point number, and networks using IF do not perform well in regression tasks (Rançon et al. 2022; Lu et al. 2021; Shrestha and Orchard 2018). It’s difficult to ensure accuracy by directly using IF for modeling. To obtain full-precision information, (Li et al. 2022) changes the Eqn. 4 to:
@@
B-FIF: Bounded full-precision integrate-and-fire spiking neuron.
@@
$u^{\text{pre}}=k\cdot r\cdot\tanh({u+Wx\over r}).$ (9)
@@
${{\partial y}\over{\partial V_{\text{th}}}}=\lambda\max(0,{k-\left|u^{\text{pre}}-V_{\text{th}}\right|\over k^{2}})u^{\text{pre}}.$ (14)
@@
We evaluate our methods in 8 scenes from the Blender dataset (Mildenhall et al. 2021) and 6 scenes from Dex-NeRF (Ichnowski et al. 2021). Following most 3D reconstruction methods (Wang et al. 2021; Wang, Skorokhodov, and Wonka 2022, 2023; Oechsle, Peng, and Geiger 2021), we also evaluate our methods in the DTU dataset (Jensen et al. 2014). These datasets cover different types of scenes or objects and are benchmarks for NeRF-based methods. Furthermore, we additionally use 2 self-created scenes to show that our method is better than previous methods in special scenarios (see supplementary material for more details).
@@
wheel tennis mount glass turbine clutter pawn pipe Avg.
@@
Lego Chair Mic Ficus Hotdog Drums Mats Ship Avg.
PET 0.58 0.65 0.59 0.71 1.02 2.53 1.05 1.57 1.09
HFS 0.96 0.65 0.72 0.87 1.35 3.82 1.08 2.18 1.45
NeuS 1.52 0.70 0.85 1.67 1.40 4.27 1.08 2.33 1.73
NeRF 2.06 0.75 0.95 0.56 1.83 3.38 1.12 3.12 1.72
Mip 1.92 0.90 1.13 0.55 1.98 3.34 1.30 2.03 1.64
Ours 0.70 0.66 0.72 0.54 0.94 2.43 1.10 1.49 1.07
@@
As shown in the 5th and 6th columns of Fig. 3, with conventional NeRF-based methods, due to the lack of surface constraints, the reconstruction results have many erroneous high-frequency information. Additionally, there is a loss of reconstruction accuracy for thin objects and semi-transparent scenes. As shown in the 2nd, 3rd, and 4th columns of Fig. 3, with SDF-based methods, the reconstruction resulted in significant errors in high-frequency changes in the image, such as failing to reconstruct holes that should have been present and surfaces that should have been separate. Moreover, NeuS demonstrated excessive smoothness in the 2nd row of Fig. 3. NeuS, HF-NeuS, and PET-NeuS fail in the reconstruction of semi-transparent scenes. Even though HF-NeuS uses some methods to reconstruct the high-frequency information, it still fails in some thin high-frequency positions. Because our method uses the hybrid ANN-SNN framework to model the 3D geometric information in a discontinuous representation, the reconstructed geometric information is more accurate.
@@
37 40 55 63 65 83 97 105 110 114 118 Avg.
PET 0.84 0.72 0.40 0.95 0.84 1.42 1.14 0.72 1.05 0.44 0.59 0.83
HFS 1.37 0.78 0.47 1.11 0.68 1.20 1.17 0.85 1.27 0.38 0.54 0.90
NeuS 1.37 0.93 0.43 1.10 0.70 1.48 1.16 0.83 1.69 0.35 0.49 0.96
NeRF 1.63 1.75 0.60 2.03 1.07 1.70 1.95 1.13 2.33 0.87 1.05 1.46
Mip 1.83 1.65 1.61 2.90 1.79 1.93 2.19 1.32 2.52 1.38 1.49 1.87
Ours 1.20 1.02 0.38 1.15 0.72 1.10 1.19 0.65 1.60 0.49 0.55 0.90
Table 3: Quantitative Comparison on DTU (Jensen et al. 2014). We show a quantitative comparison for the reconstructed geometry on 11 scans from the DTU dataset.
@@
To assess the effectiveness of the initial smoothing, we design an experiment without $L_{\text{g}}$ and compare the result with our method. It can be seen that although the Chamfer distances are similar, the reconstructed surface contains wrong high-frequency information (see the 4th column of Fig. 5), indicating that training from the wrong high-frequency surface is difficult.
@@ Conclusion
We present a novel neural surface reconstruction method for reconstructing objects. NeRF and its variants use volume rendering to produce a neural scene representation. However, extracting high-quality surfaces from this learned implicit representation is difficult because there is not a definite interface density. In our work, we propose to use a hybrid ANN-SNN framework to reconstruct density fields. We observe that the conventional volume rendering method causes inherent geometric errors for surface reconstruction, and propose a new spiking neuron to get more accurate surface reconstruction with an adaptive spiking threshold.
@@ Acknowledgements
Supplementary Material for “Spiking NeRF: Representing
the Real-World Geometry by a Discontinuous Representation”
This supplementary material consists of two parts, the additional analysis and the additional results.
Figure 6: The blue line and orange line in the right represent two different light. The figures on the left represent the distribution of the density field through which light passes. $t$ is the sampling distance from the sampling point to the camera origin. $d_{1}$ and $d_{2}$ represent $t$ corresponding to the position of surface points along the direction of light rays (i.e., the depth corresponding to the direction of light). $v_{1}$ and $v_{2}$ represent the density corresponding to the position of surface points along the direction of light rays. For continuous fields, using the same threshold to extract geometric information cannot guarantee that the depth corresponding to the light is all accurate.
Additional Analysis
Optimal Threshold Perturbation
For a well-trained NeRF, $\int_{0}^{T}{\sigma(t)\cdot\exp\left({-\int_{0}^{t}{\sigma(t)}\,{\rm d}t}\right)\cdot t}\,{\rm d}t$ can represent the depth of the corresponding light direction (Zhang et al. 2021; Guo et al. 2022). However, it depends on the direction. The final geometric information is extracted from the network using a threshold, independent of the direction. As shown in Fig. 6, when light from different directions passes through a density field, the corresponding density distribution is different, resulting in different density values corresponding to depth. So using the same threshold to extract geometric information cannot guarantee that the depth corresponding to the light is all accurate.
The Proof of Proposition 1
Define $m$ is the index of the first non zero $\sigma$ (i.e., the extracted depth $d_{\text{v}}=t_{m}$), and $m^{\prime}$ is the index of the largest $\sigma$ within m+1 and N (i.e., $V_{max}=\sigma_{m^{\prime}}$). $\Delta t_{i}$ is the sampling interval between sampling point $i-1$ and point $i$. The $w_{i}$ refers to the weight at each sampling point as defined in (Mildenhall et al. 2021). We define $\beta_{i}=e^{-\sigma_{i}\Delta t_{i}}$. Then we have:
$\sum_{i}^{N}w_{i}=1-\prod_{i}^{N}\beta_{i}.$ (15)
Let $c=\prod_{i=1}^{N}\beta_{i}$. Since $\sigma_{1}$, …, $\sigma_{m-1}$ are all 0, we have:
$0<\beta_{m}e^{-\sigma_{m^{\prime}}(t_{N}-t_{m})}<c<\beta_{m}\beta_{m^{\prime}}$ (16)
According to previous analysis, for well trained NeRF, it can be considered that the depth $d$ is $\int_{0}^{T}{\sigma(t)\cdot\exp\left({-\int_{o}^{t}{\sigma(t)}\,{\rm d}t}\right)\cdot t}\,{\rm d}t$. Its Riemann sum form is $\sum_{i=1}^{N}{w_{i}t_{i}}$.
$d-t_{m}$ $=\sum_{i=1}^{N}{w_{i}t_{i}}-t_{m}$ (17)
$=\sum_{i=m}^{N}{w_{i}t_{i}}-t_{m}\left(\sum_{i=1}^{N}{w_{i}}+c\right)$
$\geq\left(1-w_{m}-c\right)\left(t_{m+1}-t_{m}\right)-t_{m}c$
$\geq\left(t_{m+1}-t_{m}\right)\beta_{m}-t_{m+1}\beta_{m}\beta_{m^{\prime}}$
$\geq\left(\Delta t_{m+1}-Te^{-\sigma_{m^{\prime}}\Delta t_{m^{\prime}}}\right)e^{-\sigma_{m}\Delta t_{m}}.$
For a well-trained NeRF, the spiking threshold $V_{\text{th}}$ equals to the non zero minimum value of its density field. Because the value before the last layer varies continuously with respect to the spatial point, the density value of the first point that a ray encounters with a non zero density equals to the non zero minimum value of its density field. Then $V_{\text{th}}=\sigma_{m}$. So we have:
$d-t_{m}$ $\geq\left(\Delta t_{m+1}-Te^{-\sigma_{m^{\prime}}\Delta t_{m^{\prime}}}\right)e^{-\sigma_{m}\Delta t_{m}}$ (18)
$=\left(\Delta t_{m+1}-Te^{-V_{\text{max}}\Delta t_{m^{\prime}}}\right)e^{-V_{\text{th}}\Delta t_{m}}.$
$d-t_{m}$ $=\sum_{i=1}^{N}{w_{i}t_{i}}-t_{m}$ (19)
$=\sum_{i=m}^{N}{w_{i}t_{i}}-t_{m}\left(\sum_{i=1}^{N}{w_{i}}+c\right)$
$\leq\left(t_{N}-t_{m}\right)\left(1-w_{m}-c\right)-t_{m}c$
$\leq\left(t_{N}-t_{m}\right)\beta_{m}-t_{N}\beta_{m}e^{-\sigma_{m^{\prime}}\left(t_{N}-t_{m}\right)}$
$\leq\left(t_{N}-t_{N}e^{-\sigma_{m^{\prime}}t_{N}}\right)\beta_{m}$
$=T\left(1-e^{-V_{\text{max}}T}\right)e^{-V_{\text{th}}\Delta t_{m}}.$
Since we do not explore the influence of sampling point positions, we have written $\Delta t_{m}$,$\Delta t_{m+1}$ and $\Delta t_{m^{\prime}}$ in the equation as $\Delta t$ in the paper for the convenience of observation.
Then we have:
[…]
Figure 7: Self created scenes. Left: A tennis racket with a board. Right: A wheel.
For a general setting, T is greater than 100 times $\Delta t$. And $\Delta t$ is approximately $0.01$. So $\left|\Delta t-Te^{-V_{\text{max}}\Delta t}\right|<=T\left(1-e^{-V_{\text{max}}T}\right)$ for $V_{\text{max}}>4$. Therefore, for a fixed $V_{\text{th}}>4$, which is basically valid, a small $V_{\text{max}}$ can decrease the error.
Figure 8: The relationship between the upper bound and the average depth error during training. We show two more scenes from Blender dataset (Mildenhall et al. 2021). The red curve represents upper bound while the blue curve represents the average depth error during training.
Other Bounded Functions
IF spiking neuron.
The IF spiking neuron has a bound. However, according to the analysis in Preliminary, it is ultimately difficult to maintain accuracy using IF spiking neuron.
FIF with a hard bound.
$y_{t+1}=\min(o_{t+1}\cdot V_{\text{th}},B).$ (22)
Here, $B$ is the upper bound of FIF spiking neuron. Although there is a bound, increasing the bound requires additional operations (e.g., additional loss term or a strategy to increase $B$), If $B$ is not updated, the network will only output 00 when $V_{\text{th}}$ exceeds $B$.
We use $k\tanh()$ in the paper because $k$ can be updated through the original loss.
Analysis of Surrogate Gradient
The surrogate gradient in (Li et al. 2022) is:
${{\partial L}\over{\partial V_{\text{th}}}}=\sum{{{\partial L}\over{\partial y_{t}}}(1-o_{t})}.$ (23)
Due to $t=1$ in our paper. We have:
${{\partial L}\over{\partial V_{\text{th}}}}={{{\partial L}\over{\partial y}}(1-o)}.$ (24)
Since the value of $1-o$ can only be 00 or $1$, it can lead to unstable training.
Additional Results
Introduction of Our Dataset
We additionally generate our own dataset containing pathtraced images of two objects. They are all rendered from viewpoints uniformly sampled on a full sphere. We render 100 views of each scene as input and 200 for testing, all at 800 × 800 pixels (see Fig. 7).
Validation for Upper Bound of Proposition 1
We show two more scenes for the upper bound of proposition 1 in Fig. 8. It can be seen that the average depth error decreases with the upper bound and the average depth error keeps being less than upper bound during training.